# Exploring Efficient Open-Vocabulary Segmentation in the Remote Sensing

### v1

@@
Exploring Efficient Open-Vocabulary Segmentation in the Remote Sensing
Open-Vocabulary Remote Sensing Image Segmentation (OVRSIS), an emerging task that adapts Open-Vocabulary Segmentation (OVS) to the remote sensing (RS) domain, remains underexplored due to the absence of a unified evaluation benchmark and the domain gap between natural and RS images. To bridge these gaps, we first establish a standardized OVRSIS benchmark (OVRSISBench) based on widely-used RS segmentation datasets, enabling consistent evaluation across methods. Using this benchmark, we comprehensively evaluate several representative OVS/OVRSIS models and reveal their limitations when directly applied to remote sensing scenarios. Building on these insights, we propose RSKT-Seg, a novel open-vocabulary segmentation framework tailored for remote sensing. RSKT-Seg integrates three key components: (1) a Multi-Directional Cost Map Aggregation (RS-CMA) module that captures rotation-invariant visual cues by computing vision-language cosine similarities across multiple directions; (2) an Efficient Cost Map Fusion (RS-Fusion) transformer, which jointly models spatial and semantic dependencies with a lightweight dimensionality reduction strategy; and (3) a Remote Sensing Knowledge Transfer (RS-Transfer) module that injects pre-trained knowledge and facilitates domain adaptation via enhanced upsampling. Extensive experiments on the benchmark show that RSKT-Seg consistently outperforms strong OVS baselines by +3.8 mIoU and +5.9 mACC, while achieving 2× faster inference through efficient aggregation. Our code is here.
@@ Introduction
Semantic segmentation, a classic task in computer vision, aims to achieve pixel-level category prediction (Chen et al. 2017). Traditional semantic segmentation models are based on manually annotated datasets, which usually cover only a limited number of categories. Similarly, traditional remote sensing image segmentation has long focused on a fixed set of predefined categories (Kotaridis and Lazaridou 2021; Diakogiannis et al. 2020). However, as application demands grow, especially in scenarios like tracking new urban infrastructure or identifying rare geological features, the shortcomings of this limited-category-based approach have become evident. This has spurred the development of open-vocabulary remote sensing image segmentation (OVRSIS) (Ye, Zhuge, and Zhang 2025).
@@
OVRSIS, builded on the general OVS concept(Li et al. 2025; Cho et al. 2024; Xie et al. 2024b; Xu et al. 2023), is tailored to the unique characteristics of remote sensing imagery. By leveraging the power of VLMs and cross-modal learning (Radford et al. 2021; Jia et al. 2021), OVRSIS breaks the constraints of traditional training categories, enabling the segmentation of novel classes not present in the original training data. This significantly improves the adaptability and generalization ability of remote sensing image segmentation.
@@
However, unlike the recent advances in OVS (Xu et al. 2023; Radford et al. 2021; Jia et al. 2021), its extension to the remote sensing (RS) domain (OVRSIS) remains largely underexplored. A critical bottleneck hindering progress in this emerging field is the absence of a standardized evaluation benchmark. Most prior works assess their models on limited datasets or under inconsistent experimental setups, making it difficult to draw fair comparisons or systematically analyze model behavior in remote sensing scenarios. To address this issue, we construct a unified benchmark for OVRSIS, named OVRSISBench, by reformulating several widely-used remote sensing segmentation datasets under open-vocabulary settings. By leveraging established RS datasets with open-vocabulary configurations, OVRSISBench retains the domain-specific challenges of traditional RS tasks while incorporating the flexibility and generalization demands of open-vocabulary segmentation. This benchmark facilitates fair, consistent, and scalable evaluation of OVS models in realistic RS environments.
@@
Using this OVRSISBench benchmark, we comprehensively evaluate several representative OVS models and observe a significant drop in performance when these models are directly transferred from natural images to the RS domain (see Fig. fig. 1(a-b)). Furthermore, we survey recent works on OVRSIS model and evaluate them under the proposed OVRSISBench. Our experiments reveal that while some approaches demonstrate incremental improvements compared with OVS model, they often fail to achieve effective knowledge transfer to the RS domain. For example, OVRS (Cao et al. 2024), which introduces only minor modifications to OVS model (Cho et al. 2024), shows performance gains but lacks mechanisms to incorporate RS-specific priors, leading to limited adaptability. Similarly, models such as GSNet (Ye, Zhuge, and Zhang 2025), although partially incorporating RS knowledge, overlook key characteristics of RS imagery such as large-scale context, object rotation, and spectral diversity.
@@
Semantic segmentation, a crucial task in computer vision, focuses on pixel-level classification. Over the years, it has seen significant progress with various methods and models. Fully convolutional networks (FCNs) (Long, Shelhamer, and Darrell 2015) were an early milestone. As end-to-end models, they enabled direct pixel-wise predictions. Later, SegNet (Badrinarayanan, Kendall, and Cipolla 2017) and U-Net (Ronneberger, Fischer, and Brox 2015) evolved from FCNs. These encoder-decoder architectures effectively captured low-level and high-level features, gaining popularity. ResNet is widely used as a feature encoder in semantic segmentation, forming the basis of many models (Lin et al. 2018). Recently, Vision Transformers (ViT) (Dosovitskiy 2020) have brought new ideas. By using self-attention, ViT models capture long-range dependencies. Adaptations like SETR (Zheng et al. 2021) improve efficiency. New paradigms based on ViT have emerged for semantic segmentation (Li et al. 2024b; Hu et al. 2024), which requires precise pixel-level feature handling (Xie et al. 2024a; Zhu et al. 2024; Li et al. 2024a). Notable works such as MaskFormer (Cheng, Schwing, and Kirillov 2021) and Mask2Former (Cheng et al. 2022) unify pixel-level and mask classification, enhancing performance. SegFormer addresses resolution issues by removing positional encodings and presenting multi-resolution features (Xie et al. 2021). We summarize works about OVS in appendix A.
@@
OVRSIS is an emerging field in a rapid development stage, and preliminary exploratory research work has been carried out. In the process of migrating from the OVS field to the remote sensing field, OVRS (Cao et al. 2024) takes into account the basic characteristics of the remote sensing field, makes some fundamental improvements on catseg, and has achieved certain results. GSNet (Ye, Zhuge, and Zhang 2025)introduces a specialist-generalist model, which can better integrate remote sensing domain knowledge into the model. SegEarth-OV (Li et al. 2024c) proposes a train-free model, which shows better performance than the OVS model on 17 datasets. Overall, as an emerging field, in OVRSIS, few existing models can consider remote sensing features and introduce remote sensing domain knowledge for migration. Most of them only modify and improve the OVS model. In view of this, we propose the RSKT-Seg model. By considering the rotational invariance of remote sensing images and introducing remote sensing domain knowledge, this model has obtained satisfactory performance on multiple datasets.
@@
OVRSISBench follows an open-vocabulary paradigm like (Cao et al. 2024), where the model must generalize to previously unseen classes based on text descriptions (fig. 2(a)). To construct this benchmark, we adapt 8 representative remote sensing datasets: DLRSD (Chaudhuri et al. 2017), iSAID (Yao et al. 2021), Potsdam, Vaihingen, UAVid(Lyu et al. 2020), UDD5(Chen et al. 2018), LoveDA(Wang et al. 2021) and VDD(Cai et al. 2025). These datasets cover a diverse range of scenes, including urban layouts, agricultural regions, and high-resolution aerial imagery, the examples of the datasets are illustrated in the fig. 2(c). We introduce the datasets in the appendix C.
@@
As shown in fig. 2(a), we utilize DLRSD and iSAID as training sets as (Cao et al. 2024) due to their large scale and diversity, enabling robust learning of remote sensing visual patterns. For evaluation, we test across all 8 datasets to ensure the model’s generalization under varying scene distributions and resolutions. The training and evaluation splits are consistent across all methods to ensure fair comparison.
@@
To rigorously assess the open-vocabulary generalization capability of the models, particularly its performance on unseen classes. We adopt a cross-dataset transfer protocol in our OVRSISBench. In this setting, the models are trained on the DLRSD and iSAID datasets and evaluated on the separate target datasets with partially disjoint category sets. This protocol aligns with the standard open-vocabulary segmentation (OVS) setting(Xu et al. 2022; Cho et al. 2024). As shown in fig. 2(b), we further perform a statistical analysis of the cross-category distribution to substantiate the open-vocabulary nature of our setup.
@@
We first evaluated the performance of several classic OVS model on OVRSISBench, and subsequently selected recently published OVRSIS methods for evaluation. For the evaluation metrics, we adopted mean Intersection over Union (mIoU) and mean Accuracy (mACC). We introduce the Evaluation metric in the appendix D. We utilized the pre-trained ViT-B/16 (Dosovitskiy 2020) and ViT-L/14@336 (Radford et al. 2021) as the vision-language models.
@@
The overall framework of our proposed RSKT-Seg is shown in Figure fig. 3(a). We design the OVRSIS architecture with remote sensing knowledge transfer. The preparatory knowledge for training pre-trained VLMs such as CLIP has been extensively introduced in previous methods (Hu et al. 2024). The definition of OVRSIS is the same as that of OVS. We will not elaborate on it here, and researchers can refer to the corresponding sections of relevant literature (Li et al. 2025; Cho et al. 2024; Xu et al. 2023).
@@
As fig. 4(a-b), we use the CLIP image encoder to encode , yielding . To embed remote-sensing domain knowledge, we employ a pre-trained DINO encoder on a large-scale dataset (Ye, Zhuge, and Zhang 2025). Encoding with , we get .
@@
The cost map is used to characterize the correlation between visual features and text features. Building upon this, we gradually predict the segmentation map by enhancing the cost map’s spatial discriminative ability ( dimension) and class discriminative ability ( dimension). To achieve this, we have designed two dedicated modules: the Spatial Enhancement Transformer (SET) and the Class Enhancement Transformer (CET), as illustrated in fig. 3(b) and (c). This idea aligns with that presented in (Cho et al. 2024). To further strengthen the integration of remote sensing domain knowledge, we embed knowledge from the pre-trained DINO model together with the cost map. Additionally, considering the requirement for faster inference speed, we have developed two feature dimension reduction methods and a lightweight transformer architecture.
@@
RS-Transfer Upsample
Here, Proj is a feature projection module. After layers of refinement, the final prediction O is generated in the original image resolution, with shape .
Training Loss
Experiment
Experimental Analysis of RSKT-Seg
Implementation Details of RSKT-Seg
Our approach is implemented based on the PyTorch and Detectron2 (Wu et al. 2019) frameworks. The number of layers in RSKT-Upsample is configured as 2. When using the ViT-B/16 architecture as encoder, is set to the values , and when using the ViT-L/14 architecture, is set to . We use DINO ViT-B/32 pretrained in (Ye, Zhuge, and Zhang 2025) and is set to . for RemoteCLIP ViT-B/16 is set to . AdamW is utilized as the optimizer, with an initial learning rate of and a weight decay of . The batch size is fixed at 8. All the experiments are carried out on 4 Nvidia 4090 GPUs, each having 24GB of memory. Our experimental results is that the optimal number of layers is 5 when using DLRSD as the training set, and 2 when using iSAID.
Results of the Proposed Method on Different Datasets
Ablation Study On RSKT-Seg
The Effectiveness of Different Components
Dataset | R-I Cost Map | DINO Cost Map | RS-Transfer | RS-Fusion | m-mIoU | m-fwIoU | m-mACC |
@@
✓ | ✓ | ✓ | ✓ | 39.80 | 41.59 | 62.55 |
The importance of introducing remote sensing knowledge
Dataset | rotate | rsDINO | natureDINO | remoteclip | m-mIoU | m-fwIoU | m-mACC |
DLRSD | 47.17 | 52.26 | 60.48 | ||||
✓ | 48.22 | 54.86 | 62.41 | ||||
✓ | 46.46 | 53.87 | 62.19 | ||||
✓ | ✓ | 49.47 | 55.45 | 62.81 | |||
✓ | ✓ | 46.90 | 54.96 | 62.40 | |||
✓ | ✓ | ✓ | 49.37 | 55.29 | 62.74 | ||
✓ | ✓ | ✓ | 47.17 | 54.11 | 61.70 |
Ablation of Different cost map fusion strategy
@@
DLRSD m-mIoU | 46.21 | 46.87 | 49.25 | 48.84 | 51.58 | 48.20 |
DLRSD m-mACC | 61.35 | 62.36 | 64.06 | 63.11 | 65.41 | 64.11 |
iSAID m-mIoU | 38.13 | 39.09 | 38.93 | 37.22 | 37.84 | 37.34 |
iSAID m-mACC | 61.32 | 62.21 | 62.47 | 60.17 | 60.29 | 60.24 |
This subsection analyzes the impacts of different layers in the cost map aggregation module using data from table 5 on DLRSD and iSAID datasets, with metrics like m-mIoU, m-fwIoU, and m-mACC. Based on the experimental results, we selected N=5 for the DLRSD dataset and N=2 for the iSAID dataset as the optimal configurations for the Cost Aggregation Module
The impacts of different finetuning strategies
@@
As seen in fig. 5, cost maps from CLIP Cost and DINO Cost show how models assign costs to classes like impervious surfaces, trees, and buildings. These heat-like visualizations reveal areas of high or low classification confidence. RS-Fusion, depicted vertically in fig. 5 across different classes, refines these maps, enhancing segmentation accuracy. More Cost visualization figures are shown in appendix E.
@@
As depicted in fig. 6, a comparison is made among the segmentation visualizations of multiple models. These models were trained on the DLRSD dataset and then tested on the Vaihingen dataset. Each row corresponds to a different model, while the columns represent various image samples. RSKT-Seg demonstrates notable effectiveness in this comparison. Compared to other models, RSKT-Seg shows a relatively lower rate of misclassification and a more precise identification of object boundaries.

### v2

@@
Abstract
Open-Vocabulary Remote Sensing Image Segmentation (OVRSIS), an emerging task that adapts Open-Vocabulary Segmentation (OVS) to the remote sensing (RS) domain, remains underexplored due to the absence of a unified evaluation benchmark and the domain gap between natural and RS images. To bridge these gaps, we first establish a standardized OVRSIS benchmark (OVRSISBench) based on widely-used RS segmentation datasets, enabling consistent evaluation across methods. Using this benchmark, we comprehensively evaluate several representative OVS/OVRSIS models and reveal their limitations when directly applied to remote sensing scenarios. Building on these insights, we propose RSKT-Seg, a novel open-vocabulary segmentation framework tailored for remote sensing. RSKT-Seg integrates three key components: (1) a Multi-Directional Cost Map Aggregation (RS-CMA) module that captures rotation-invariant visual cues by computing vision-language cosine similarities across multiple directions; (2) an Efficient Cost Map Fusion (RS-Fusion) transformer, which jointly models spatial and semantic dependencies with a lightweight dimensionality reduction strategy; and (3) a Remote Sensing Knowledge Transfer (RS-Transfer) module that injects pre-trained knowledge and facilitates domain adaptation via enhanced upsampling. Extensive experiments on the benchmark show that RSKT-Seg consistently outperforms strong OVS baselines by +3.8 mIoU and +5.9 mACC, while achieving 2× faster inference through efficient aggregation.
Code — https://github.com/LiBingyu01/RSKT-Seg
Extended version — https://arxiv.org/pdf/2509.12040
@@ Introduction
Semantic segmentation, a classic task in computer vision, aims to achieve pixel-level category prediction (chen2017deeplab). Traditional semantic segmentation models are based on manually annotated datasets, which usually cover only a limited number of categories. Similarly, traditional remote sensing image segmentation has long focused on a fixed set of predefined categories (kotaridis2021remote; diakogiannis2020resunet). However, as application demands grow, especially in scenarios like tracking new urban infrastructure or identifying rare geological features, the shortcomings of this limited-category-based approach have become evident. This has spurred the development of open-vocabulary remote sensing image segmentation (OVRSIS) (ye2025GSNet).
@@
OVRSIS, builded on the general OVS concept(li2025fgaseg; cho2024cat; xie2024sed; xu2023side), is tailored to the unique characteristics of remote sensing imagery. By leveraging the power of VLMs and cross-modal learning (radford2021learning; jia2021scaling), OVRSIS breaks the constraints of traditional training categories, enabling the segmentation of new classes not presented in the original datasets. This significantly improves the adaptability and generalization ability of remote sensing image segmentation.
@@
However, unlike the recent advances in OVS (xu2023side; radford2021learning; jia2021scaling), its extension to the remote sensing (RS) domain (OVRSIS) remains largely underexplored. A critical bottleneck hindering progress in this emerging field is the absence of a standardized evaluation benchmark. Most prior works assess their models on limited datasets or under inconsistent experimental setups, making it difficult to draw fair comparisons or systematically analyze model behavior in remote sensing scenarios. To address this issue, we construct a unified benchmark for OVRSIS, named OVRSISBench, by reformulating several widely-used remote sensing segmentation datasets under open-vocabulary settings. By leveraging established RS datasets with open-vocabulary configurations, OVRSISBench retains the domain-specific challenges of traditional RS tasks while incorporating the flexibility and generalization demands of open-vocabulary segmentation. This benchmark facilitates fair, consistent, and scalable evaluation of OVS models in realistic RS environments.
@@
Using this OVRSISBench benchmark, we comprehensively evaluate several representative OVS models and observe a significant drop in performance when these models are directly transferred from natural images to the RS domain (see Fig. fig. 1(a-b)). Furthermore, we survey recent works on OVRSIS model and evaluate them under the proposed OVRSISBench. Our experiments reveal that while some approaches demonstrate incremental improvements compared with OVS model, they often fail to achieve effective knowledge transfer to the RS domain. For example, OVRS (cao2024open), which introduces only minor modifications to OVS model (cho2024cat), shows performance gains but lacks mechanisms to incorporate RS-specific priors, leading to limited adaptability. Similarly, models such as GSNet (ye2025GSNet), although partially incorporating RS knowledge, overlook key characteristics of RS imagery such as large-scale context, object rotation, and spectral diversity.
@@
Semantic segmentation, a crucial task in computer vision, focuses on pixel-level classification. Over the years, it has seen significant progress with various methods and models(10887773). Fully convolutional networks (FCNs) (long2015fully) were an early milestone. As end-to-end models, they enabled direct pixel-wise predictions. Later, SegNet (badrinarayanan2017segnet) and U-Net (ronneberger2015u; wang2024towards) evolved from FCNs. These encoder-decoder architectures effectively captured low-level and high-level features, gaining popularity. ResNet is widely used as a feature encoder in semantic segmentation, forming the basis of many models (lin2018deeptongue; li2025toward; li2024unionformer). Recently, Vision Transformers (ViT) (dosovitskiy2020image; WuLZZBZR24; WuZLCJ25) have brought new ideas. By using self-attention, ViT models capture long-range dependencies. Adaptations like SETR (zheng2021rethinking) improve efficiency. New paradigms based on ViT have emerged for semantic segmentation (li2024u3m; hu2024contrastive; yu2025mildmultilayerdiffusionstrategy), which requires precise pixel-level feature handling (xie2024multi; zhu2024saswot; li2024stitchfusion). Notable works such as MaskFormer (cheng2021per) and Mask2Former (cheng2022masked) unify pixel-level and mask classification, enhancing performance. SegFormer addresses resolution issues by removing positional encodings and presenting multi-resolution features (xie2021segformer). We summarize works about OVS in appendix A.
@@
OVRSIS is an emerging field in a rapid development stage, and preliminary exploratory research work has been carried out. In the process of migrating from the OVS field to the remote sensing field, OVRS (cao2024open) takes into account the basic characteristics of the remote sensing field, makes some fundamental improvements on catseg, and has achieved certain results. GSNet (ye2025GSNet)introduces a specialist-generalist model, which can better integrate remote sensing domain knowledge into the model. SegEarth-OV (li2024segearth) proposes a train-free model, which shows better performance than the OVS model on 17 datasets. Overall, as an emerging field, in OVRSIS, few existing models can consider remote sensing features and introduce remote sensing domain knowledge for migration. Most of them only modify and improve the OVS model. In view of this, we propose the RSKT-Seg model. By considering the rotational invariance of remote sensing images and introducing remote sensing domain knowledge, this model has obtained satisfactory performance on multiple datasets.
@@
OVRSISBench follows an open-vocabulary paradigm like (cao2024open), where the model must generalize to previously unseen classes based on text descriptions (fig. 2(a)). To construct this benchmark, we adapt 8 representative remote sensing datasets: DLRSD (chaudhuri2017multilabel), iSAID (yao2021scale), Potsdam, Vaihingen, UAVid(LYU2020108), UDD5(chen2018large), LoveDA(NEURIPS_DATASETS_AND_BENCHMARKS2021_4e732ced) and VDD(cai2025vdd). These datasets cover a diverse range of scenes, including urban layouts, agricultural regions, and high-resolution aerial imagery, the examples of the datasets are illustrated in the fig. 2(c). We introduce the datasets in the appendix C.
@@
As shown in fig. 2(a), we utilize DLRSD and iSAID as training sets as (cao2024open) due to their large scale and diversity, enabling robust learning of remote sensing visual patterns. For evaluation, we test across all 8 datasets to ensure the model’s generalization under varying scene distributions and resolutions. The training and evaluation splits are consistent across all methods to ensure fair comparison.
@@
To rigorously assess the open-vocabulary generalization capability of the models, particularly its performance on unseen classes. We adopt a cross-dataset transfer protocol in our OVRSISBench. In this setting, the models are trained on the DLRSD and iSAID datasets and evaluated on the separate target datasets with partially disjoint category sets. This protocol aligns with the standard open-vocabulary segmentation (OVS) setting(xu2022simple; cho2024cat). As shown in fig. 2(b), we further perform a statistical analysis of the cross-category distribution to substantiate the open-vocabulary nature of our setup.
@@
We first evaluated the performance of several classic OVS model on OVRSISBench, and subsequently selected recently published OVRSIS methods for evaluation. For the evaluation metrics, we adopted mean Intersection over Union (mIoU) and mean Accuracy (mACC). We introduce the Evaluation metric in the appendix D. We utilized the pre-trained ViT-B/16 (dosovitskiy2020image) and ViT-L/14@336 (radford2021learning) as the vision-language models.
@@
The overall framework of our proposed RSKT-Seg is shown in Figure fig. 3(a). We design the OVRSIS architecture with remote sensing knowledge transfer. The preparatory knowledge for training pre-trained VLMs such as CLIP has been extensively introduced in previous methods (hu2024contrastive). The definition of OVRSIS is the same as that of OVS. We will not elaborate on it here, and researchers can refer to the corresponding sections of relevant literature (li2025fgaseg; cho2024cat; xu2023side).
@@
As fig. 4(a-b), we use the CLIP image encoder to encode , yielding . To embed remote-sensing domain knowledge, we employ a pre-trained DINO encoder on a large-scale dataset (ye2025GSNet). Encoding with , we get .
@@
The cost map is used to characterize the correlation between visual features and text features. Building upon this, we gradually predict the segmentation map by enhancing the cost map’s spatial discriminative ability ( dimension) and class discriminative ability ( dimension). To achieve this, we have designed two dedicated modules: the Spatial Enhancement Transformer (SET) and the Class Enhancement Transformer (CET), as illustrated in fig. 3(b) and (c). This idea aligns with that presented in (cho2024cat). To further strengthen the integration of remote sensing domain knowledge, we embed knowledge from the pre-trained DINO model together with the cost map. Additionally, considering the requirement for faster inference speed, we have developed two feature dimension reduction methods and a lightweight transformer architecture.
@@
| Dataset | Proposed Components | Metrics | |||||
| R-I | DINO | RST | RSF | mIoU | fwIoU | mAcc | |
@@
| ✓ | ✓ | ✓ | ✓ | 39.80 | 41.59 | 62.55 | |
RS-Transfer Upsample
Here, Proj() is a feature projection module. After layers of refinement, the final prediction O is generated in the original image resolution, with shape .
| Pretrained / Augmented Models | Metrics | |||||
| rotate | RS-DINO | N-DINO | RCLIP | mIoU | fwIoU | mAcc |
| 47.17 | 52.26 | 60.48 | ||||
| ✓ | 48.22 | 54.86 | 62.41 | |||
| ✓ | 46.46 | 53.87 | 62.19 | |||
| ✓ | ✓ | 49.47 | 55.45 | 62.81 | ||
| ✓ | ✓ | 46.90 | 54.96 | 62.40 | ||
| ✓ | ✓ | ✓ | 49.37 | 55.29 | 62.74 | |
| ✓ | ✓ | ✓ | 47.17 | 54.11 | 61.70 | |
Training Loss
@@
| DLRSD | 46.21 | 46.87 | 49.25 | 48.84 | 51.58 | 48.20 |
| DLRSD | 61.35 | 62.36 | 64.06 | 63.11 | 65.41 | 64.11 |
| iSAID | 38.13 | 39.09 | 38.93 | 37.22 | 37.84 | 37.34 |
| iSAID | 61.32 | 62.21 | 62.47 | 60.17 | 60.29 | 60.24 |
Experiment
Experimental Analysis of RSKT-Seg
Implementation Details of RSKT-Seg
Our approach is implemented based on the PyTorch and Detectron2 (wu2019detectron2) frameworks. The number of layers in RSKT-Upsample is configured as 2. When using the ViT-B/16 architecture as encoder, is set to the values , and when using the ViT-L/14 architecture, is set to . We use DINO ViT-B/32 pretrained in (ye2025GSNet) and is set to . for RemoteCLIP ViT-B/16 is set to . AdamW is utilized as the optimizer, with an initial learning rate of and a weight decay of . The batch size is fixed at 8. All the experiments are carried out on 4 Nvidia 4090 GPUs, each having 24GB of memory. Our experimental results is that the optimal number of layers is 5 when using DLRSD as the training set, and 2 when using iSAID.
@@
Results of the Proposed Method on Different Datasets
Ablation Study On RSKT-Seg
The Effectiveness of Different Components
The importance of introducing remote sensing knowledge
Ablation of Different cost map fusion strategy
The impacts of different layers of cost map aggregation
This subsection analyzes the impacts of different layers in the cost map aggregation module using data from table 5 on DLRSD and iSAID datasets, with metrics like m-mIoU, m-fwIoU, and m-mACC. Based on the experimental results, we selected N=5 for the DLRSD dataset and N=2 for the iSAID dataset as the optimal configurations for the Cost Aggregation Module.
The impacts of different finetuning strategies
The analysis of inference speed
@@
As seen in fig. 5, cost maps from CLIP Cost and DINO Cost show how models assign costs to different classes. These cost reveal areas of high or low classification confidence. RS-Fusion, depicted vertically across different classes, refines these maps, enhancing segmentation accuracy. More Cost visualizations are shown in appendix E.
@@
As depicted in fig. 6, a comparison is made among the segmentation visualizations of multiple models. These models were trained on the DLRSD dataset and then tested on the Vaihingen dataset. Each row corresponds to a different model, while the columns represent various image samples. Compared to other models, RSKT-Seg shows a more precise identification of object boundaries.